OpenStudy (anonymous):

Which property does 3(17) = 3(20) – 3(3) illustrate?

OpenStudy (dape):

It illustrates the distributive property of multiplication, or the fact that \[ a(b+c)=a*b+a*c \] In this case \[ 3(17)=3(20-3)=3*20-3*3 \]